OpenStudy (anonymous):

Use finite sum to estimate the average value of the function on the given interval by partitioning the interval and evaluating the function at the midpoints of the sub intervals f(x)=3x^5 on [1,3] divided into 4 subintervals

OpenStudy (anonymous):

so what are the subintervals? 1-1.5, 1.5-2, 2-2.5, 2.5-3 midpoints?

OpenStudy (anonymous):

I'm gonna be honest this problem is greek to me....

OpenStudy (anonymous):

well, it's a way of estimating the area under the curve. do you see that? know what it's about?

OpenStudy (anonymous):

actually, here the average value of the function we divide the interval into pieces, pick some points, compute f(x) at each point, and divide by the number of points.

OpenStudy (anonymous):

Ok so that makes sense.....

OpenStudy (anonymous):

So random points in each interval?

OpenStudy (anonymous):

No, read the problem again.

OpenStudy (anonymous):

mid points

OpenStudy (anonymous):

of subintervals? 1-1.5, 1.5-2, 2-2.5, 2.5-3

OpenStudy (anonymous):

How do I find those

OpenStudy (anonymous):

midpoint of 1-1.5 = 1.25 midpoint of 1.5 -2 = ?

OpenStudy (anonymous):

oh I get it so now I just substitute those values for x?

OpenStudy (anonymous):

1.5-2=1.75

OpenStudy (anonymous):

we divide the interval into pieces, pick some points, compute f(x) at each point, and divide by the number of points.

OpenStudy (anonymous):

the points being 1.25, 1.75, 2.25, 2.75

OpenStudy (anonymous):

so something like 3(1.25)^5+3(1.75)^5+...

OpenStudy (anonymous):

I'm not very good at arithmetic :)

OpenStudy (anonymous):

ok but is my thought process correct? then I just divide by the number of subintervals (4)

OpenStudy (anonymous):

so it would be 3(1.25)^5+3(1.75)^5+3(2.25)^5+3(2.75)^5=703.22 703.22/4=175.8

OpenStudy (anonymous):

Yes